# U-PINet: Physics-Informed Hierarchical Learning for Accurate and Fast 3D RCS Prediction

### v2

@@ V-B2 Generalization
V-C 3D Object Reconstruction Time Results
Table III summarizes the reconstruction time in seconds of Fig. 6’s cube, sphere, and cone objects, as well as Fig. 7’s SLICY and ship targets, using the U-PINet, the conventional EM scattering modeling approaches MoM and MLFMA, as well as MLFMA-PINN and ConvNet. It can be observed that the U-PINet achieves a substantial speed of computation advantage over physics-based solvers, reducing runtime reductions of several orders of magnitude across both canonical (cube, sphere, and cone) and complex (SLICY and ship) 3D geometries. For example, the U-PINet completes inference on the ship model in only $0.12$ seconds, whereas the MoM requires over $80000$ seconds for the same task. Compared to the MLFMA-PINN, the U-PINet reduces runtime by up to three orders of magnitude, while preserving competitive accuracy. It is finally shown that the ConvNet runs slightly faster than the proposed model, however, ConvNet’s error rates and instability under geometric variation makes it unsuitable for practical EM applications. Concluding, the results in this table underscore U-PINet’s capability as a fast and reliable surrogate model for EM simulations.
@@
V-D Ablation Capability
We further investigate the effect of the physics-constrained loss on the error distribution of U-PINet during the current induction process. The reconstruction error distributions under the physics-informed loss in (24) and the purely data-driven MSE loss are compared for both the training and testing stages on the slender forebody model in Fig. 7 (a), as shown in Fig. 15.
@@ V-D Ablation Capability
To quantify the contributions of the physics-aware and interpretable modules in U-PINet, an ablation study is conducted on the complex ship target shown in Fig. 7(c). The quantitative results, evaluated using the four reconstruction metrics in Section IV-B, are summarized in Table IV. Specifically, the full U-PINet is compared with two degraded variants, namely, a model without the physics-informed loss in (24) and another with the edge attention module removed. As shown in Table IV, removing either component consistently degrades the regression performance.
Specifically, when the physics-informed loss is excluded, the model reverts to a purely data-driven approach, causing the RMSE to deteriorate significantly from 2.65{2.65} dB to $3.45$ dB, while the MAE increases from 2.29{2.29} dB to $2.95$ dB. This substantial performance drop underscores the critical role of physical constraints in regularizing the solution space and mitigating overfitting on limited data. Similarly, removing the edge attention module results in a noticeable accuracy loss, with the RMSE rising to $2.93$ dB and the MAPE increasing from 0.1750{0.1750} to $0.2450$. This degradation confirms that the attention mechanism is essential for capturing high-frequency scattering features localized at sharp geometric edges, which are dominant in the RCS of complex targets like the ship. Ultimately, the full U-PINet achieves the lowest errors and the highest correlation ($R^{2}=0.8950$), demonstrating that the synergistic integration of physical laws and attention-based feature extraction is key to high-fidelity RCS prediction.
These results indicate that the physics-informed loss primarily enforces global physical consistency and suppresses large scattering errors, while the edge attention module refines local behavior around geometric discontinuities. Their combination is therefore essential for achieving the full U-PINet performance.
VI Conclusion and Future Directions
In this paper, we proposed U-PINet, an end-to-end physics-informed framework for efficient 33D RCS prediction. By synergizing a near-field graph module with a hierarchical far-field network, U-PINet eliminates the need for iterative solvers while strictly adhering to electromagnetic principles. Experimental evaluations on complex targets, such as the SLICY and ship models, demonstrate that the proposed method achieves solver-level RCS accuracy and strong generalization to unseen geometries, all while reducing computational time by orders of magnitude compared to traditional MoM and MLFMA solvers.

### v4

@@ VI-B2 Generalization
VI-C 3D Object Reconstruction Time Results
Table III summarizes the reconstruction time in seconds of Fig. 6’s cube, sphere, and cone objects, as well as Fig. 7’s SLICY and ship targets, using the U-PINet, the conventional EM scattering reconstruction approaches MoM and MLFMA, as well as MLFMA-PINN and ConvNet. It can be observed that the U-PINet achieves a substantial speed of computation advantage over physics-based solvers, reducing runtime reductions of several orders of magnitude across both canonical (cube, sphere, and cone) and complex (SLICY and ship) 3D geometries. For example, the U-PINet completes inference on the ship model in only $0.12$ seconds, whereas the MoM requires over $80000$ seconds for the same task. Compared to the MLFMA-PINN, the U-PINet reduces runtime by up to three orders of magnitude, while preserving competitive accuracy. It is finally shown that the ConvNet runs slightly faster than the proposed model, however, ConvNet’s error rates and instability under geometric variation makes it unsuitable for practical EM applications. Concluding, the results in this table underscore U-PINet’s capability as a fast and reliable surrogate model for EM simulations.
@@
VI-D Ablation Capability
We further investigate the effect of the physics-informed residual loss on the error distribution of U-PINet during the current induction process. The reconstruction error distributions under the physics-informed loss in (24) and the purely data-driven MSE loss are compared for both the training and testing stages on the slender forebody model in Fig. 7 (a), as shown in Fig. 16.
@@ VI-D Ablation Capability
To quantify the contributions of the physics-aware and interpretable modules in U-PINet, an ablation study is conducted on the complex ship target shown in Fig. 7(c). The quantitative results, evaluated using the four reconstruction metrics in Section V-B, are summarized in Table IV. Specifically, the full U-PINet is compared with two degraded variants, namely, a model without the physics-informed loss in (24) and another with the edge attention module removed. As shown in Table IV, removing either component consistently degrades the regression performance.
Specifically, when the physics-informed loss is excluded, the model reverts to a purely data-driven approach, causing the RMSE to deteriorate significantly from 2.65{2.65} dB to $3.45$ dB, while the MAE increases from 2.29{2.29} dB to $2.95$ dB. This substantial performance drop underscores the critical role of physical constraints in regularizing the solution space and mitigating overfitting on limited data. Similarly, removing the edge attention module results in a noticeable accuracy loss, with the RMSE rising to $2.93$ dB and the MAPE increasing from 0.1750{0.1750} to $0.2450$. This degradation confirms that the attention mechanism is essential for capturing high-frequency scattering features localized at sharp geometric edges, which are dominant in the RCS of complex targets like the ship. Ultimately, the full U-PINet achieves the lowest errors and the highest correlation ($R^{2}=0.8950$), demonstrating that the synergistic integration of physical laws and attention-based feature extraction is key to high-fidelity RCS prediction. These results indicate that the physics-informed loss primarily enforces global physical consistency suppressing large scattering errors, while the edge attention module refines local behavior around geometric discontinuities. Their combination is therefore essential for achieving the full U-PINet performance.
VII Conclusion and Future Directions
In this paper, we propose U-PINet, an end-to-end physics-informed framework for efficient RCS prediction via 3D EM scattering reconstruction. The proposed artificial neural network reconstructs the fundamental scattering quantity of the induced surface currents, and then evaluates the scattered field and RCS through the radiation integral and far-field definition, rather than directly regressing RCS. By combining a physics-guided near-field graph module for local coupling with a hierarchical far-field module for long-range radiation, U-PINet attains solver-level RCS accuracy on both canonical and complex targets with orders of magnitude runtime reduction, while generalizing well to unseen geometries and varying observation conditions under limited training data.